OpenStudy (anonymous):

What is the simplified form of ^3 square root x^2• ^3 square root x^2

OpenStudy (anonymous):

what are the choices

jhonyy9 (jhonyy9):

i think these wan being cuberot x^2 * cuberoot x^2 = ? but not is sure

OpenStudy (kittiwitti1):

my interpretation:\[\sqrt[3]{x^{2}\times\sqrt[3]{x^2}}\]

OpenStudy (kittiwitti1):

yes I think you are right jhony

jhonyy9 (jhonyy9):

ok. but the asker can decid it sure what wan asking here truly

OpenStudy (anonymous):

Jhonyy you're right

OpenStudy (anonymous):

The options are: X^3 sqareroot x ^3 square root x X X^4/9

OpenStudy (kittiwitti1):

@jhonyy9

jhonyy9 (jhonyy9):

yes i m here

jhonyy9 (jhonyy9):

ok so you can rewrite them like exponential form (x^2)^(1/3) *(x^2)^(1/3) = ? this will help you ?

OpenStudy (anonymous):

I think? Is it x^4/9